# Topic Modeling for Short Texts with Large Language Models

### v1

@@ C.1 Examples of Topics
C.2 Examples of Topic Words Not Included in the Documents
Table 12 shows examples of words not included in the documents outputted in topic modeling on GoogleNewsT. GPT-3.5 output in bold are the names of entities (e.g., broncos, gree, and watson) or words that do not exist in the real world (e.g., dorffiefskee).
Such words are considered harmful because they may induce misinterpretation of topics. However, only a small number of such words were found, and most of them were synonyms, synonyms, derivatives, or related words in the documents.
C.3 Examples of the Processing
Table C.3 reports the concrete examples of topics identified for each subset and the final output to demonstrate the processing in our methods. In the parallel approach, we find that LLM reasonably merges topics from each subset. For instance, bold topics in each subset are merged into one topic in the final output, using words from both subsets. On the other hand, in the sequential approach, the final output is the same as the topics for the first subset except for only one pair of bold words. This indicates that LLMs with the the sequential approach could too strictly keep topics for the previous subset so that they can not output topics that sufficiently reflect the entire set.
Appendix D Limitations
We do not consider it matters whether GPT-3.5 and GPT-4 have seen these datasets in the pre-training; since topic modeling is an unsupervised task, they can not utilize their knowledge about these datasets in our experiment. Our approaches perform promisingly in a practical common situation where the number of topics is about 10 but are less reliable with a larger number of topics, such as 50 or 100, due to the difficulty of controlling the LLM’s output format.
@@ Appendix D Limitations
| Model | Examples of the topic words not included in the documents |
| TACTM | twelvemonth sink railway blowout |

### v2

@@ C.1 Examples of Topics
C.2 Analysis of Parallel and Sequential Prompting
Table 2 and Table 3 show that the parallel prompting approach can identify topics with better coherence and document coverage than the sequential prompting one. To analyze the superior performance of the parallel approach, we calculate Cv and DC of topics before merging. Table 12 shows that Cv and DC scores before merging degrade from those of the parallel approach, demonstrating that the merging process can improve both their coherence and document coverage.
Additionally, we examine the transition of topics during their processing and then observe that the sequential approach tends to update the previously identified topic very little due to strict adherence to our instruction, leading to lower document coverage compared to the parallel approach. Table 13 reports the concrete examples of topics identified for each subset and the final output to demonstrate the processing in our methods. In the parallel approach, we find that LLM reasonably merges topics from each subset. For instance, bold topics in each subset are merged into one topic in the final output, using words from both subsets. In contrast, in the sequential approach, the final output is the same as the topics for the first subset except for only one pair of bold words. This indicates that LLMs with the the sequential approach could too strictly keep topics for the previous subset so that they can not output topics that sufficiently reflect the entire set.
C.3 Examples of Topic Words Not Included in the Documents
Table 14 shows examples of words not included in the documents outputted in topic modeling on GoogleNewsT. GPT-3.5 output in bold are the names of entities (e.g., broncos, gree, and watson) or words that do not exist in the real world (e.g., dorffiefskee). Such words are considered harmful because they may induce misinterpretation of topics. However, only a small number of such words were found, and most of them were synonyms, synonyms, derivatives, or related words in the documents.
Appendix D Controllability of the Number of Topics
Unlike conventional topic models that consistently produce the specified number of topics, LLMs may not generate the exact number of topics indicated in the prompt. Therefore, to evaluate LLM’s controllability of the number of topics, we investigate its output when the specified number of topics in prompts is set to 5, 10, 15, and 20, while keeping the number of topic words fixed at 5. We use the documents set from Section 5 and GPT-4 (gpt-4-turbo-2024-04-09) with Base prompt, running ten trials for each setting. The results show that the same number of topics as specified are outputted in all trials, demonstrating GPT-4’s sufficient controllability of the number of topics.
@@ Appendix D Controllability of the Number of Topics
| Model size | Examples of Llama2 Output |
| # of | Text | Vocabulary | |
| Dataset | Documents | Length | Size |
| Tweet | 2000 | 5.47 | 706 |
| GoogleNewsT | 11000 | 5.25 | 2376 |
| StackOverFlow | 19000 | 4.71 | 2544 |
| 20NG (all samples) | 500 | 147.25 | 2095 |
| 20NG (computer) | 126 | 98.27 | 1437 |
| 20NG (politics) | 77 | 296.69 | 1792 |
| 20NG (science) | 57 | 113.51 | 1672 |
@@
| Model | Tweet | GoogleNewsT | StackOverFlow | |||||
| Cv | DC | Cv | DC | Cv | DC | |||
| GPT-3.5 | 0.532 | 0.366 | 0.517 | 0.569 | 0.464 | 0.634 | ||
| GPT-4 | 0.580 | 0.395 | 0.523 | 0.665 | 0.519 | 0.747 | ||
@@
| Model | Examples of the topic words not included in the documents |
| TACTM | twelvemonth sink railway blowout |